# Classical bounds on correlation-type Bell expressions and linear prepare-and-measure witnesses: efficient computation in parallel environments such as graphics processing units

### v1

@@ Abstract
The presented program aims at speeding up the brute force computation of the $L_{d}$ norm of a matrix $M$ using graphics processing units (GPUs). Alternatives for CPUs have also been implemented, and the algorithm is applicable to any parallel environment. The $n\times m$ matrix $M$ has real elements which may represent coefficients of a bipartite correlation-type Bell expression or those of a linear prepare-and-measure (PM) witness. In this interpretation, the $L_{1}$ norm is the local bound of the given Bell expression, and the $L_{d}$ norm for $d\geq 2$ is the classical $d$-dimensional bound of the given PM witness, which is associated with the communication of $d$-level classical messages in the PM scenario. In both scenarios, the output is assumed to be binary.
The code for GPUs is written in CUDA C and can utilize one NVIDIA GPU in a computer. To illustrate the performance of our implementation, we refer to Brierley et al. [1] who needed approximately three weeks to compute the local bound on a Bell expression defined by a $42\times 42$ matrix on a standard desktop using a single CPU core. In contrast, our efficient implementation of the brute force algorithm allows us to reduce this to three minutes using a single NVIDIA RTX 6000 Ada graphics card on a desktop. For CPUs, the algorithm was implemented with OpenMP and MPI according to the shared and distributed memory models, respectively, and achieves a comparable speedup at a number of CPU cores around 100.
keywords:
journal: Computer Physics Communications
PROGRAM SUMMARY/NEW VERSION PROGRAM SUMMARY
Program Title: L_CUDA.cu, L_MPI.c, L_OpenMP.c
@@
Supplementary material:
Journal reference of previous version:*
Does the new version supersede the previous version?:*
Reasons for the new version:*
Summary of revisions:
@@ 1 Introduction
In this paper, we present efficient brute force computations of both the local bound of correlation-type Bell expressions and the one-dit classical bound of linear witnesses in the prepare-and-measure scenario using parallel environments such as graphics processing units (GPUs). The local (classical) bound of a generic correlation-type Bell expression, irrespective of the dimensionality, is given by [1, 21, 22]
@@ 2.1 $L_{1}$ norm
Consider a matrix $M$ with $n$ rows and $m$ columns. Consider a strategy vector $\mathbf{a}$ of dimension $n$, where the vector entries can be $+1$ or $-1$ resulting in $2^{n}$ possible options for the vector $\mathbf{a}$. The $L_{1}$ norm of $M$ was defined in Eq. (1) through all those strategy vectors as the optimal value among the Manhattan norms of the vectors $\mathbf{m}=\mathbf{a}M$. Computing the $L_{1}$ norm naively would thus mean $2^{n}$ matrix multiplications with a vector. We have identified four ways to reduce the number of operations to calculate the $L_{1}$ value. Two of them correspond to a prior transformation of the matrix, while the other two concern the actual algorithm representing the brute-force computation. We will discuss them all below, before describing parallelization.
@@ 2.1.2 Algorithmic speedups
An immediate algorithmic simplification arises from the fact that the $L^{*}_{1}$ Bell values are the same for a given strategy $\mathbf{a}$ and for the opposite strategy vector, $-\mathbf{a}$. Then all possible $L^{*}_{1}$ Bell values will remain covered if we fix one entry of the strategy vector. The number of strategy vectors to be considered for calculating all possible $L^{*}_{1}$ Bell values thereby decreases to $2^{n-1}$. (Note that this already implies that the optimal strategy vector is not unique as the optimal $L^{*}_{1}$ Bell value can be attained by two different strategy vectors at least; but even more strategy vectors may be optimal in other cases.)
@@
To describe this method we need to consider a series of vectors having a Hamming distance of one like the column vectors in Table 1. A simple choice is the so-called binary reflected Gray code (BRGC) [30, 31]; see Eqs. (16)-(17) for a definition in case $d=2$, and refer to Table 4 for a recursive and more pictorial formulation. Note that the BRGC entries $G$ take a value of 00 or $1$ according to the definition; however, conversion to entries $a=1$ or $a=-1$ is straightforward through $a=2G-1$, which results in a one-to-one correspondence between the strategy vectors and the columns of the BRGC.
@@
$j$th word
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
$i$th digit 3 0 0 0 0 0 0 0 0 1 1 1 1 1 1 1 1
2 0 0 0 0 1 1 1 1 1 1 1 1 0 0 0 0
1 0 0 1 1 1 1 0 0 0 0 1 1 1 1 0 0
0 0 1 1 0 0 1 1 0 0 1 1 0 0 1 1 0
3- 0+ 1+ 0- 2+ 0+ 1- 0- 3+ 0+ 1+ 0- 2- 0+ 1- 0-
Table 1: The four-digit binary reflected Gray code (BRGC). The columns show words of the BRGC. The Hamming distance of the neighboring vectors is 1. The last row shows where the $j$th vector differs from the previous one and also shows in the right upper index if it is modified from zero to one ($+$) or from one to zero (--). (For $j=0$, comparison is made with the last word.) The index $j$ of the word or vector of the BRGC and $i$ of the digit of the BRGC are shown in the upper row and left column, respectively.
row index $i$
3 3 3
2 2 3 2
1 1 2 2 2 1
0 0 1 1 1 1 1 1 1 0
@@ 2.1.3 Parallelization and further technical aspects
Considering the way the algorithm determines the $j$th word of the Gray code, by Eqs. (12) and (13), it gives a constraint on the length of the Gray code the program can handle. This constraint is determined by the number of bits the integer is represented on a computer. If the integer is represented on $B$ bits, the program can determine Gray codes of length $B-2$ having $2^{B-2}$ words. As the variables are defined as unsigned long long int which is stored in 64 bits, the code can calculate Gray codes of length 62. As the code considers one row with a fixed sign, the strategy vector can be of length $B-1$, which is 63 for 64-bit unsigned integers. Since the program transposes the input matrix if the number of columns exceeds the number of rows, the eventual constraint is that the number of rows and the number of columns of the input matrix cannot be more than 63 simultaneously.
@@ 4 Final remarks
In this article, we have presented a method to efficiently calculate the $L_{d}$ norm of a matrix, for any given $d\geq 1$, in a parallel environment. Apart from their significance in the evaluation of the classical bounds on Bell inequalities or prepare-and-measure witnesses, these norms are also of interest in the field of communication complexity, the Grothendieck constant, or in graph theory. As a result of parallelization (with special emphasis on GPUs) and the use of special mathematical and programming techniques, our implementation has achieved a significant speedup factor in comparison with previous computations performed on a single CPU core (see e.g. related numerics-focused works under Refs. [34, 35, 1, 36] on computing the $L_{1}$ norm). We believe that our primitives can also be used as an improved simulation of the oracle in the optimization algorithms of Refs. [1, 37, 38, 39], thus further boosting their performance.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper….
@@ Acknowledgments
We thank Tamás Vértesi for valuable input and fruitful discussions, and Csaba Hruska for providing technical assistance in the GPU computations. This paper was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. We acknowledge the support of the EU (CHIST-ERA MoDIC) and the National Research, Development and Innovation Office NKFIH (No. 2023-1.2.1-ERA_NET-2023-00009 and No. K145927).
@@
Note that an arbitrary column can be multiplied by $-1$ without affecting the $L_{1}$ norm, which ensures consistency with the freedom to choose an arbitrary order for subtraction. Since transposing the matrix leaves $L_{1}$ unchanged, our proof also covers the case of rows besides that of columns.
@@ Appendix B Proving the invariance of $L_{d}$, $d\geq 2$, when adding specific columns
The $(n+1)$-digit case is generated recursively from the $n$-digit case by reflecting the $n\times 2^{n}$ matrix $G$ and completing the first row with values 00 for column indices $10\leq j\leq 2^{n}-$ 1 and with values $1$ for $12^{n}\leq j\leq 2^{n+1}-$ 1, as it is depicted in Table 4.
@@
For $12^{n}\leq j\leq 2^{n+1}-$ 1, the definition and assumption (33) imply
@@
According to Eq. (12), the $j$th and $(j-1)$th column can be expressed as
@@
If $R=0$, the expressions $\left\lfloor Z+\frac{R}{2^{i+1}}\right\rfloor$ and $\left\lfloor Z+\frac{R}{2^{i+1}}-\frac{1}{2^{i+1}}\right\rfloor$ have different parities as a result of the inequality $0<\frac{1}{2^{i+1}}\leq\frac{1}{2}$. If $R\neq 0$, these expressions have the same parities, because $\frac{R}{2^{i+1}}<1$ and $10\leq\frac{R}{2^{i+1}}-\frac{1}{2^{i+1}}<$ 1. This means that the change occurs when the following condition is satisfied:
@@ Appendix E Proving that the changes in the $d$-ary Gray code groups occur at the same digits
According to Eq. (18), the change occurs at the digit with the maximal $i$ such that $j$ is divisible by $d^{i}$. Since $k\neq 0$, $j$ is not divisible by $d^{h-l}$. This means $j$ is not divisible by $d^{i}$ if $i\geq h-l$. This implies $i<h-l$ for the digit $i$ of the change.

### v2

# Classical bounds on two-outcome bipartite Bell expressions and linear prepare-and-measure witnesses: Efficient computation in parallel environments such as graphics processing units

@@ Abstract
The presented program aims at speeding up the brute force computation of the so-called $L_{d}$ norm of a matrix $M$ using graphics processing units (GPUs). Alternatives for CPUs have also been implemented, and the algorithm is applicable to any parallel environment. The $n\times m$ matrix $M$ has real elements which may represent coefficients of a bipartite Bell expression or those of a linear prepare-and-measure (PM) witness. In this interpretation, the $L_{1}$ norm is the local bound of the given correlation-type Bell expression, and the $L_{d}$ norm for $d\geq 2$ is the classical $d$-dimensional bound of the given PM witness, which is associated with the communication of $d$-level classical messages in the PM scenario. The program is also capable of calculating the local bound of Bell expressions including marginals. In all scenarios, the output is assumed to be binary.
The code for GPUs is written in CUDA C and can utilize one NVIDIA GPU in a computer. To illustrate the performance of our implementation, we refer to Brierley et al. [1] who needed approximately three weeks to compute the local bound on a Bell expression defined by a $42\times 42$ matrix on a standard desktop using a single CPU core. In contrast, our efficient implementation of the brute force algorithm allows us to reduce this to three minutes using a single NVIDIA RTX 6000 Ada graphics card on a workstation. For CPUs, the algorithm was implemented with OpenMP and MPI according to the shared and distributed memory models, respectively, and achieves a comparable speedup at a number of CPU cores around 100.
PROGRAM SUMMARY Program Title: L_CUDA.cu, L_MPI.c, L_OpenMP.c
@@ 1 Introduction
In this paper, we present efficient brute force computations of both the local bound of Bell expressions and the one-dit classical bound of linear witnesses in the prepare-and-measure scenario using parallel environments such as graphics processing units (GPUs). The local (classical) bound of a generic correlation-type Bell expression, irrespective of the dimensionality, is given by [1, 21, 22]
@@
In a generic Bell expression Alices’s (Bob’s) marginal coefficients are defined by the first row (column) of the Bell matrix [28, 29, 30]. It is easy to show that the local bound can be expressed in this case as
$L_{marg}(M):=\max_{\left\{\begin{subarray}{c}a_{1}=+1\\ a_{x}=\pm 1,\text{ }x\geq 2\end{subarray}\right\}}\left(\sum_{x=1}^{n}M_{x1}a_{x}+\sum_{y=2}^{m}\left|\sum_{x=1}^{n}M_{xy}a_{x}\right|\right).$ (2)
@@ 2.1 $L_{1}$ norm
Consider a matrix $M$ with $n$ rows and $m$ columns. Consider a strategy vector $\mathbf{a}$ of dimension $n$, where the vector entries can be +1+1 or 1-1 resulting in $2^{n}$ possible options for the vector $\mathbf{a}$. The $L_{1}$ norm of $M$ was defined in Eq. (1) through all those strategy vectors as the optimal value among the Manhattan norms of the vectors $\mathbf{m}=\mathbf{a}M$. Computing the $L_{1}$ norm naively would thus mean $2^{n}$ matrix multiplications with a vector. We have identified four ways to reduce the number of operations to calculate the $L_{1}$ value. Two of them correspond to a prior transformation of the matrix, while the other two concern the actual algorithm representing the brute-force computation. We will discuss them all below, before describing parallelization.
@@ 2.1.2 Algorithmic speedups
An immediate algorithmic simplification arises from the fact that the $L^{*}_{1}$ Bell values are the same for a given strategy $\mathbf{a}$ and for the opposite strategy vector, $\mathbf{a}$. Then all possible $L^{*}_{1}$ Bell values will remain covered if we fix one entry of the strategy vector. The number of strategy vectors to be considered for calculating all possible $L^{*}_{1}$ Bell values thereby decreases to $2^{n-1}$. (Note that this already implies that the optimal strategy vector is not unique as the optimal $L^{*}_{1}$ Bell value can be attained by two different strategy vectors at least; but even more strategy vectors may be optimal in other cases.)
@@
To describe this method we need to consider a series of vectors having a Hamming distance of one like the column vectors in Table 1. A simple choice is the so-called binary reflected Gray code (BRGC) [33, 34]; see Eqs. (17)-(18) for a definition in case $d=2$, and refer to Table 4 for a recursive and more pictorial formulation. Note that the BRGC entries $G$ take a value of 0 or $1$ according to the definition; however, conversion to entries $a=1$ or a=1a=-1 is straightforward through a=2G1a=2G-1, which results in a one-to-one correspondence between the strategy vectors and the columns of the BRGC.
@@
$j$th word
0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15
$i$th digit
3
0 0 0 0 0 0 0 0 1 1 1 1 1 1 1 1
2 0 0 0 0 1 1 1 1 1 1 1 1 0 0 0 0
1 0 0 1 1 1 1 0 0 0 0 1 1 1 1 0 0
0 0 1 1 0 0 1 1 0 0 1 1 0 0 1 1 0
3- 0+ 1+ 0- 2+ 0+ 1- 0- 3+ 0+ 1+ 0- 2- 0+ 1- 0-
Table 1: The four-digit binary reflected Gray code (BRGC). The columns show words of the BRGC. The Hamming distance of the neighboring vectors is 1. The last row shows where the $j$th vector differs from the previous one and also shows in the right upper index if it is modified from zero to one (++) or from one to zero (-). (For $j=0$, comparison is made with the last word.) The index $j$ of the word or vector of the BRGC and $i$ of the digit of the BRGC are shown in the upper row and left column, respectively.
row index $i$
3 3 3
2 2 3 2
1 1 2 2 2 1
0 0 1 1 1 1 1 1 1 0
@@ 2.1.3 Parallelization and further technical aspects
Considering the way the algorithm determines the $j$th word of the Gray code, by Eqs. (13) and (14), it gives a constraint on the length of the Gray code the program can handle. This constraint is determined by the number of bits the integer is represented on a computer. If the integer is represented on $B$ bits, the program can determine Gray codes of length B2B-2 having $2^{B-2}$ words. As the variables are defined as unsigned long long int which is stored in 64 bits, the code can calculate Gray codes of length 62. As the code considers one row with a fixed sign, the strategy vector can be of length B1B-1, which is 63 for 64-bit unsigned integers. Since the program transposes the input matrix if the number of columns exceeds the number of rows, the eventual constraint is that the number of rows and the number of columns of the input matrix cannot be more than 63 simultaneously.
2.2 Special notes for calculating $L_{marg}$
In case one intends to calculate the local bound of Bell expressions including marginals (see Eq. 2), the previous statements for preprocessing the witness matrix remain valid with the following constraints:
  • 1.
    Trivially, when all entries of a given row or column are zero except for the first row or column; such a row or column can be removed. The first row or column can be removed when both the entries of the first row and the entries of the first column are exclusively zero. In this case the program will calculate the $L_{1}$ norm of the reduced matrix.
  • 2.
    When a row or column of the matrix is a constant multiple of another row or column except for the first row or column. In this case, the two rows or columns can be replaced by a single one by adding or subtracting the two, depending on the sign of the constant multiplier.
According to Eq. 2, the rest of the computation can be performed as described above except that the fixed entry of the strategy vector must be chosen to be the first one and the first column needs to be treated separately from the rest.
@@ 2.3.3 Technical considerations
3 A simple analysis of the speedup
As we are claiming that our speedup relies on several factors, it is natural to elaborate on the issue and perform some test for verification. Parallelization itself is an obvious source of speedup, and we will illustrate here that our solution provides more both in theory and practice. We will concentrate on the $L_{1}$ norm of $n\times n$ square matrices as an example, and our reference will be the algorithm used by Brierley et al. [1] which is perhaps the most naive implementation of computing this norm.
In particular, a brute force computation must consider approximately $2^{n}$ different strategy vectors, which is the main contribution in the scaling of the computation time with $n$. For each strategy vector, Brierley et al. computed the Bell value from “scratch”, which requires performing $n^{2}$ operations (a matrix-matrix multiplication in their code). However, an iterative algorithm like ours does this at the beginning only, and updates the Bell value for each new strategy vector by accessing a single row of the matrix in question, resulting in $n$ operations in each such iteration. Hence, the naive implementation is expected to scale with $n$ as $n^{2}2^{n}$, whereas an iterative one as $n2^{n}$.
Our particular code uses BRGC to achieve an iterative implementation. In order to compare its performance with the algorithm used by Brierley et al., we implemented the latter in C to eliminate the difference in performance between programming languages. As we are interested here in the algorithmic speedup as opposed to parallelization, we ran our programs utilizing a single thread; concretely, on an Intel(R) Xeon(R) Gold 6448Y processor after compiling it by the GNU Compiler Collection with no optimization (-O0). We simply measured the elapsed time for the two methods with matrices of different sizes; to reduce random contributions from unrelated system load, we actually performed the measurement three times and took the minimum.
Figure 2: Execution times for computing the $L_{1}$ norm for matrices with different sizes ($n\times n$) using a single core of an Intel(R) Xeon(R) Gold 6448Y processor. The codes were written in C and compiled using the GNU Compiler Collection with no optimization (-O0). The blue dots depict the execution times for the code using our algorithm, meanwhile, the red dots depict the execution times for a naive code (see text). Least-squares fits are indicated in the plot.
The measurement results can be seen in Fig. 2. The red and blue dots are depicting the execution times as the function of the size $n$ of the matrices using the algorithm of Brierly et al. and ours, respectively. Functions of the form $f(n)=A\times n^{2}\times 2^{n}$ and $g(n)=B\times n\times 2^{n}$, respectively, fit very well to the measurement results, where the constants $A$ and $B$ have been fitted according to least squares. This confirms our advantage of a factor of $n$ in terms of scaling; furthermore, the prefactor $B$ turns out to be similar to $A$, indicating that our algorithm indeed has a major algorithmic advantage for matrices of relevant sizes.
Our results and some further experience (not shown) also suggest that administrative costs at the beginning or during the bulk of the computation are negligible in comparison with the full execution time. In fact, this is why we did not include an additive constant into functions $f(n)$ and $g(n)$. As an example for administrative costs during the bulk, calculating the next strategy vector does have a computational cost. However, in the case of Brierley et al., the calculation of the strategy vector out of a non-negative number requires bit operations, which is very efficient. On the other hand, our iterative algorithm determines the index where the next BRGC word has changed. One can notice from Eq. 13 and Table 1 that $1/2$ of the time the change between the consecutive BRGC words occur at the 0th digit, $1/4$ of the time it happens at the 1st digit, $1/8$ of the time it happens at the 2nd digit and so on. If we consider an arbitrarily chosen word of an $n$-digit BRGC, the expected value of the times we need to calculate Eq. 14 is $\sum_{i=1}^{n}\frac{i}{2^{i}}$. As $\sum_{i=1}^{\infty}\frac{i}{2^{i}}=2$ and $n<\infty$, we conclude that computing the next BRGC word in our algorithm is not a task that could strongly influence execution times and their scaling with $n$. We also need to mention that Eq. 14 is implemented with bit operations as well.
@@ 5 Final remarks
In this article, we have presented a method to efficiently calculate the $L_{d}$ norm of a matrix, for any given $d\geq 1$, in a parallel environment. Apart from their significance in the evaluation of the classical bounds on Bell inequalities or prepare-and-measure witnesses, these norms are also of interest in the field of communication complexity, the Grothendieck constant, or in graph theory. In order to be able to consider all two-outcome bipartite Bell inequalities, our program can also calculate the local bound of Bell expressions including marginals. As a result of parallelization (with special emphasis on GPUs) and the use of special mathematical and programming techniques, our implementation has achieved a significant speedup factor in comparison with previous computations performed on a single CPU core (see e.g. related numerics-focused works under Refs. [37, 38, 1, 39] on computing the $L_{1}$ norm). We believe that our primitives can also be used as an improved simulation of the oracle in the optimization algorithms of Refs. [1, 40, 41, 42], thus further boosting their performance.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
We thank Tamás Vértesi, Kai-Siang Chen and Hsin-Yu Hsu for valuable input and fruitful discussions. The authors thank Csaba Hruska for providing technical assistance in the GPU computations. This paper was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences, as well as the EU (CHIST-ERA MoDIC) and the National Research, Development and Innovation Office NKFIH (No. 2023-1.2.1-ERA_NET-2023-00009 and No. K145927). G.D. acknowledges support from the EU (European Social Fund Plus) and the Government of the Balearic Islands through a Vicenç Mut postdoctoral fellowship (No. PD-035-2023).
@@
Note that an arbitrary column can be multiplied by 1-1 without affecting the $L_{1}$ norm, which ensures consistency with the freedom to choose an arbitrary order for subtraction. Since transposing the matrix leaves $L_{1}$ unchanged, our proof also covers the case of rows besides that of columns.
@@ Appendix B Proving the invariance of $L_{d}$, $d\geq 2$, when adding specific columns
The $(n+1)$-digit case is generated recursively from the $n$-digit case by reflecting the $n\times 2^{n}$ matrix $G$ and completing the first row with values 0 for column indices $0\leq j\leq 2^{n}-1$ and with values $1$ for $2^{n}\leq j\leq 2^{n+1}-1$, as it is depicted in Table 4.
@@
For $2^{n}\leq j\leq 2^{n+1}-1$, the definition and assumption (34) imply
@@
According to Eq. (13), the $j$th and (j1)(j-1)th column can be expressed as
@@
If $R=0$, the expressions $\left\lfloor Z+\frac{R}{2^{i+1}}\right\rfloor$ and $\left\lfloor Z+\frac{R}{2^{i+1}}-\frac{1}{2^{i+1}}\right\rfloor$ have different parities as a result of the inequality $0<\frac{1}{2^{i+1}}\leq\frac{1}{2}$. If $R\neq 0$, these expressions have the same parities, because $\frac{R}{2^{i+1}}<1$ and $0\leq\frac{R}{2^{i+1}}-\frac{1}{2^{i+1}}<1$. This means that the change occurs when the following condition is satisfied:
@@ Appendix E Proving that the changes in the $d$-ary Gray code groups occur at the same digits
According to Eq. (19), the change occurs at the digit with the maximal $i$ such that $j$ is divisible by $d^{i}$. Since $k\neq 0$, $j$ is not divisible by $d^{h-l}$. This means $j$ is not divisible by $d^{i}$ if $i\geq h-l$. This implies i<hli<h-l for the digit $i$ of the change.